# CADDesigner: Conceptual Design of CAD Models Based on General-Purpose Agent

### v2

@@ Introduction
Multi-view snapshots are taken of the generated CAD model $\mathcal{M}$, resulting in an image set $\mathcal{V}=\{Img_{1},Img_{2},\dots,Img_{6}\}$ (corresponding to the front top left, back bottom right, back top left, front bottom right, top and right views respectively). CADDesigner uses the function
$Check:\mathcal{D}\times\mathcal{V}\to\{0,1\}$
@@
to check whether the model meets the user’s intent. When $Check(\mathcal{D},\mathcal{V})=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $Check(\mathcal{D},\mathcal{V})=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, $Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage. In the entire process, except for the user feedback links ($R$ and SatSat), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of StateState are independently completed by the CADDesigner).
CAD Code Generation Paradigm
@@
In this section, we provide a detailed description of a new CAD code generation paradigm from four key aspects: the LLM-friendly Context-Independent Imperative Paradigm (CIP), structured tracking and error information, and the automatic construction of a reusable modeling knowledge base.
Definition
We define two mappings: one between the language system and the semantic space, and the other between two language systems.
Definition 1 (Language System, Express Mapping and Explain Mapping):
Formally, language system is defined as a tuple $\mathcal{L}=(\Sigma^{*},\mathbf{R})$.
where $\Sigma^{*}$ denotes the set of all finite strings over a finite alphabet. $\mathcal{M}$ is the semantic space.
  • Express Mapping $\mathbf{R}$ that converts strings to semantic elements is defined as:
    $\mathbf{R}:\Sigma^{*}\rightarrow\mathcal{M}$
    And,
    $\forall s\in\Sigma^{*},\quad\exists\mathcal{M}^{\prime}\subsetneq\mathcal{M}\quad\text{s.t.}\quad\mathbf{R}(s)\subseteq\mathcal{M}^{\prime}$
  • Explain Mapping $\mathbf{R}^{-1}$ that recovers strings from semantic elements is defined as:
    $\mathbf{R}^{-1}:\mathcal{M}\rightharpoonup\Sigma^{*}$
    This mapping is partial, meaning it is not necessarily defined for all elements in $\mathcal{M}$. Specifically,
    $\exists m\in\mathcal{M}\quad\text{s.t.}\quad\mathbf{R}^{-1}(m)=\varnothing,$
    indicating that certain semantics cannot be explained or reconstructed using strings in the given language system.
Definition 2 (Translation Mapping):
Let $\mathcal{L}_{A}=(\Sigma_{A}^{*},\mathbf{R}_{A})$ and $\mathcal{L}_{B}=(\Sigma_{B}^{*},\mathbf{R}_{B})$ be two language systems over the same semantic space $\mathcal{M}$. The translation mapping from $\Sigma_{A}^{*}$ to $\Sigma_{B}^{*}$ is defined as the composition:
$\mathbf{T}_{A\rightarrow B}:=\mathbf{R}_{B}^{-1}\circ\mathbf{R}_{A},$
where for $s_{A}\in\Sigma_{A}^{*}$, the translation $s_{B}=\mathbf{T}_{A\rightarrow B}(s_{A})$ is the result of first mapping $s_{A}$ into its semantic meaning in $\mathcal{M}$ via $\mathbf{R}_{A}$, then generating a corresponding string $s_{B}\in\Sigma_{B}^{*}$ using $\mathbf{R}_{B}^{-1}$.
Context-Independent Imperative Paradigm
To address the limitations of chaining modeling language such as CadQuery, we propose a new modeling paradigm termed CIP (Context-Independent Imperative Paradigm). We give the following definitions based on the definition above.
Definition 3 (CAD Operation Semantic Space):
The CAD operation semantic space is defined as a set of tuples representing morphisms:
$\forall m$ $\in\mathcal{M}_{\text{CADOP}}$
$m$ $:=(\mathcal{F}_{\textbf{OP}},[\textbf{Obj}],\textbf{Params})$
Here $\mathcal{F}_{\textbf{OP}}$ is a morphism defined as:
$\mathcal{F}_{\textbf{OP}}(\mathcal{O})=\mathcal{O}^{\prime},\quad\text{where }\mathcal{O},\mathcal{O}^{\prime}\in\textbf{Obj}\cup\textbf{Params}$
Obj denotes instances of geometric entities such as Solid, Face, etc., and Params refers to parameters required for a specific CAD operation.
Definition 4 (Context-Independent Imperative Paradigm (CIP)):
We define CIP as a language system:
$\mathcal{L}_{\textbf{CIP}}$ $=(\Sigma^{*}_{\textbf{CIP}},\mathbf{R}_{\textbf{CIP}})$
where:
  • $\forall s\in\Sigma^{*}_{\textbf{CIP}},\quad s:=\texttt{Operation}([\texttt{Obj}],\texttt{Params}),$ $s$ is a string or called a ’statement’,
  • $\forall s\in\Sigma^{*}_{\textbf{CIP}},\ \exists!m\in\mathcal{M}_{\text{CADOP}}\quad\text{s.t.}\quad\mathbf{R}_{\textbf{CIP}}(s):=m=(\mathcal{F}_{\mathbf{OP}},[\textbf{Obj}],\textbf{Params})$
E.g., $s=\texttt{extrude\_rsolid(S, direction, distance)}$, in which extrude_rsolid is an Operation, S is an Object and $\texttt{direction},\texttt{distance}$ are Parameters.
Each statement in CIP forms a self-contained and semantically closed unit. The mapping $\mathbf{R}_{\textbf{CIP}}$ yields a unique operational triple without requiring any execution history or context. This ensures that CIP is a truly context-independent paradigm.
LLM Friendly Error Information
CadQuery’s error messages are typically raw Python exceptions with unstructured stack traces. Due to its reliance on method chaining and implicit states, errors often occur far from their actual cause, providing little semantic guidance. As a result, LLMs struggle to identify how to fix the underlying issue.
In CIP, we introduce fine-grained error handling nearly at the level of each atomic operation. Each failure triggers a structured diagnostic message that includes likely causes and suggested fixes which are derived from patterns observed in common failure cases. This design ensures that errors carry actionable signals, making it much easier for LLMs to locate and repair faults.
Formally, each error is represented as:
$\mathit{ErrMsg}=(\mathit{ErrCau},\ \mathit{ErrLoc},\ \mathit{CorrAct}),$
where $\mathit{ErrCau}$ denotes the root cause of failure, $\mathit{ErrLoc}$ indicates the precise location in the user script, and $\mathit{CorrAct}$ suggests one or more potential corrective actions.
This format enables LLMs to directly extract relevant debugging signals. In addition, corrective hints are programmatically injected based on the error type and recent execution trace, offering high-level semantic suggestions rather than raw exceptions. Since all runtime logs and structured diagnostics are centrally managed, the LLM agent can access and reason over them during iterative code generation.
Explicit Type Annotation
A common source of failure observed in PythonOCC and CadQuery stems from LLMs misinterpreting the return types of certain methods. This is particularly problematic in distinguishing semantically similar but hierarchically distinct geometric entities, such as Wire and Edge. While both may appear interchangeable in conventional geometry contexts, they represent different levels of abstraction in CAD systems, which is a distinction critical for correct modeling behavior.
To address this, CIP emphasizes type clarity through both API naming conventions and type-annotated documentation. Specifically, each API function in CIP follows the pattern:
ActionName_rReturnType
@@
Here, ActionName denotes the CAD operation, r is a fixed prefix indicating “returns”, and ReturnType explicitly declares the type of the returned object. This ensures that even without inspecting documentation, the function signature itself conveys type information, reducing ambiguity during LLM inference.
In addition, each function is accompanied by detailed DocStrings with explicit type annotations, which not only guide human developers but also enhance LLM performance in code-generation tasks by reinforcing accurate type semantics during prompt-based retrieval or few-shot learning.
These design choices are later validated through ablation studies, demonstrating their effectiveness in reducing type-related errors and improving modeling success rates.
Self-Evolving by Composition Operation
Our system supports a comprehensive set of atomic CAD operations, which includes extrusion, filleting, chamfering, sweeping, lofting, etc., covering nearly the entire design operation space. Beyond this foundation, we allow the explicit definition and preservation of composite operations formed by sequences of atomic ones during the modeling process. These composites can be treated as new atomic operations in subsequent tasks, effectively turning hard-to-discover modeling strategies into reusable knowledge components and expanding the agent’s action space with higher-level, user-defined primitives. Base on Definition 4, we can easily get the Corollary that composition operation is also context-independent and bijective. We observed that the agent could directly invoke preserved composite operations to generate parts such as screws and flanges, which significantly improved the success rate of one-shot model generation.
Code Framwork Knowedge Base Construction
To enhance modeling accuracy and reduce repetitive agent failures, we construct a structured knowledge base $\mathcal{K}$ that serves as external memory during code generation. The knowledge base consists of two key components:
@@
  • Function Annotations ($\mathit{Anno}$): semantic descriptions of commands, including parameter formats, return types, and usage notes;
  • Case Examples ($\mathit{Case}$): verified code snippets preserved as reference patterns.
@@
To automate the construction of the knowledge base $\mathcal{K}$, we design a rule-based pipeline that extracts and organizes information directly from the source code. Specifically, we collect function signatures, parameter types, and return types from Python annotations and type hints. These are then aligned with the corresponding natural language descriptions found in the function-level docstrings.
@@
Experiments and Comparison
We evaluate our CAD modeling code generation algorithm on a workstation equipped with an workstation with an 8-core Intel CPU, using Python 3.12 as the runtime environment. The agent system was implemented on a customed framework tailored to our task, and retrieval-augmented knowledge base access is provided by RAGFlow.
As for LLM API service, we use Claude-4-Sonnet as main agent, Gemini-2.5-Pro for code generator, and Gemini-2.5-Flash for refining requirements.
Evaluation Metrics
To evaluate the geometric fidelity of the generated CAD models, we adopt several standard shape similarity metrics, including IoU (Intersection over Union), CD (Chamfer Distance), and HD (Hausdorff Distance). These metrics assess how closely the generated shapes match the ground-truth geometry from different perspectives.
In addition, we introduce three process-oriented metrics to analyze modeling stability and reliability under ablation settings. Pass@1 denotes the proportion of cases where valid code is produced on the first attempt, reflecting generation precision. AVG Re (Average Retry) captures the average number of retries required during the iterative correction loop, indicating the ease of convergence. SUC refers to the percentage of successful cases.
Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models, with $\mathcal{X}$ and $\mathcal{Y}$ representing their corresponding point clouds, whose mathematical formulations are described as follows:
Intersection over Union (IoU) is utilized to measure the similarity between generated models and the ground truth.
$\mathrm{IoU}(\mathcal{G},\mathcal{S})=\frac{\mathcal{G}\cap\mathcal{S}}{\mathcal{G}\cup\mathcal{S}},$
where $\mathcal{G}\cap\mathcal{S}$ denotes the overlapping volume between the reference and generated models, and $\mathcal{G}\cup\mathcal{S}$ represents their combined volumetric union. A value of 11 indicates perfect alignment, while 0 signifies no overlap.
Chamfer Distance (CD) calculates point-wise proximity between point clouds sampled from $\mathcal{X}$ and $\mathcal{Y}$:
$\mathrm{CD}(\mathcal{X},\mathcal{Y})=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\min_{y\in\mathcal{Y}}\|x-y\|_{2}^{2}+\frac{1}{|\mathcal{Y}|}\sum_{y\in\mathcal{Y}}\min_{x\in\mathcal{X}}\|y-x\|_{2}^{2}.$
Hausdorff distance(HD) is a metric for quantifying the similarity between two point sets, primarily used for evaluating the resemblance of object contours.
$\mathrm{HD}(\mathcal{X},\mathcal{Y})=\max\left\{\sup_{x\in\mathcal{X}}\inf_{y\in\mathcal{Y}}d(x,y),\sup_{y\in\mathcal{Y}}\inf_{x\in\mathcal{X}}d(x,y)\right\}$
where, $d(x,y)$ is the Euclidean distance between points $x$ and $y$, $\sup$ (supremum) and $\inf$ (infimum) represent the least upper bound and greatest lower bound, respectively.
Ablation Study on CIP Paradigm
@@
To illustrate the benefits of certain design choices in CIP, such as the error handling mechanism that is friendly to large language models (LLMs) and the overall architecture of the CIP system itself, we conducted an ablation study on Text2CAD dataset. Specifically, we evaluated four variants on the same dataset: CadQuery, CIP, CIP without the detailed error system, and CIP without the return type annotations. The results are presented in Table 1.
@@
Table 1: Ablation study of CIP components and comparative evaluation with CadQuery.
Method Pass@1↑ AVG Re↓ SUC↑
CIP 0.44 1.88 100%
CIP w/o Err 0.44 2.67 77.8%
CIP w/o Type 0.33 2.33 88.9%
CadQuery 0.18 4.00 72.7%
The results show that CIP achieves the best overall performance, clearly outperforming CadQuery in all metrics. When the error handling mechanism is removed (CIP w/o Err), the Pass@1 remains at 0.44, but the AVG Re increases from 1.88 to 2.67, and SUC drops to 77.8%. This suggests that while the agent can still generate correct code initially, the lack of detailed feedback slows down convergence during iterative correction.
@@
In contrast, removing type annotations (CIP w/o Type) causes Pass@1 to drop significantly to 0.33, indicating that the initial code generation is more prone to semantic errors without explicit type guidance. Interestingly, AVG Re (2.33) is lower than in the error-free setting but still higher than full CIP, and SUC also drops to 88.9%. This shows that type annotations help encode critical modeling intent into the prompt, but some type errors still get corrected through retries.
Overall, we observe a clear division of roles: error handling helps the agent quickly recover from mistakes during iterations, and type annotations improve first-attempt accuracy by reducing semantic ambiguity.
Comparison Methods
Figure 3: Comparison of generation results across Text2CAD, CADCodeVerify, cadrille, and our method. CADDesigner achieves the best input alignment.

### v4

@@ 3.5 Visual-based Iterative Error Correction
Multi-view snapshots of the generated CAD model $\mathcal{M}$ are taken by $T_{4}^{1}$, producing an image set $\mathcal{V}=\{\text{Img}_{1},\allowbreak\text{Img}_{2},\allowbreak\dots,\allowbreak\text{Img}_{6}\}$ (corresponding to the Front-Top-Left, Back-Bottom-Right, Back-Top-Left, Front-Bottom-Right, Top, and Right views respectively). CADDesigner evaluates whether the model meets the user’s intent $\mathcal{D}_{\text{final}}$ using a binary variable $P\in\{0,1\}$. Internally, $P$ is determined by $T_{4}^{2}$, which involves two components:
  • Visual Question Generation $F_{\mathrm{vq}}:\mathcal{V}\times\mathcal{D}_{\text{final}}\to\mathcal{Q}$, which produces a set of targeted visual questions $\mathcal{Q}=\{q_{1},q_{2},\ldots,q_{n}\}$ based on the user’s requirements and the multi-view images $\mathcal{V}$.
@@
  • Visual Feedback Generation $F_{\mathrm{vf}}:\mathcal{V}\times\mathcal{Q}\to(P,F)$, which analyzes the multi-view images $\mathcal{V}$ to answer the questions in $\mathcal{Q}$, and outputs explicit visual feedback $F$ along with the binary decision $P$. When $P=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $P=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, $Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage.
In the entire process, except for the user feedback links ($R$ and $Sat$), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of $State$ are independently completed by the CADDesigner).
4 CAD Code Generation Paradigm
@@
In this section, we provide a detailed description of a new CAD code generation paradigm from three key aspects: the LLM-friendly Explicit Context Imperative Paradigm (ECIP), LLM-friendly CAD code design, and the automatic construction of a reusable modeling knowledge base.
4.1 Explicit Context Imperative Paradigm
In recent years, many works employ CadQuery (CQ) code as an intermediate representation for CAD model generation. It adopts a Fluent API design, where operations are expressed via chained method calls to enable a highly readable and compact modeling workflow. This style relies on an implicit internal context $C$ that tracks the current modeling state and is automatically passed between operations. Formally, each operation $f_{k}$ updates the hidden state $C_{k}$ with parameters $p_{k}$:
$C_{k+1}=f_{k}(C_{k},p_{k}),$ (4)
and a chained user expression corresponds to function composition:
$C_{n}=f_{n}\circ f_{n-1}\circ\cdots\circ f_{1}(C_{0}).$ (5)
While concise, this implicit state obscures semantic dependencies, complicating understanding and automated code generation—especially for Large Language Models that struggle to track hidden contexts. In contrast, low-level APIs like PythonOCC require explicit inputs for each operation, yielding clear semantics but verbose, fine-grained code that lacks high-level abstraction and usability.
To address these limitations, we introduce the Explicit Context Imperative Paradigm (ECIP), which enforces explicit passing of all input objects and parameters to each operation. Formally, each operation $g_{k}$ maps explicit input state $S_{k}$ and parameters $p_{k}$ to a new state $S_{k+1}$:
$S_{k+1}=g_{k}(S_{k},p_{k}).$ (6)
Thus, the modeling workflow is represented as a sequence of explicit state transformations:
$S_{1}\!=\!g_{1}(S_{0},p_{1}),\,S_{2}\!=\!g_{2}(S_{1},p_{2}),\,\ldots,\,S_{n}\!=\!g_{n}(S_{n-1},p_{n}),$ (7)
where all intermediate states $S_{k}$ are explicit variables.
Figure 3: Code comparison between CadQuery (left) and ECIP (right). ECIP explicitly passes context and supports standard Python constructs, improving code clarity and flexibility.
Figure 3 illustrates a typical code comparison between CadQuery and ECIP. CadQuery (left) uses chained method calls that implicitly pass context, making the code compact but less flexible for inserting standard Python statements or loops. In contrast, ECIP (right) explicitly passes all inputs, allowing the use of standard Python control flow like $\mathtt{for}$ loops and variable assignments, which enhances modularity, readability, and ease of debugging. Thus, ECIP combines the clarity of low-level APIs with the expressiveness of high-level modeling, greatly improving the accuracy and interpretability of agent-driven CAD code generation.
4.2 LLM-Friendly CAD Code Design
In addition to the core Explicit Context Imperative Paradigm (ECIP) described in Section 4.1, we incorporate several complementary design features aimed at improving Large Language Model (LLM) agents’ ability to accurately generate and debug CAD modeling code.
4.2.1 Structured Error Information
The error handling mechanism of CadQuery typically produces generic or unstructured error messages, which significantly impede automated debugging and code correction. To address this limitation, ECIP introduces fine-grained structured error messages at the atomic operation level, represented as triples:
$\mathtt{ErrMsg}=(\mathtt{ErrCau},\ \mathtt{ErrLoc},\ \mathtt{CorrAct}),$ (8)
where $\mathtt{ErrCau}$ identifies the root cause of the failure, $\mathtt{ErrLoc}$ specifies the exact location of the error in the script, and $\mathtt{CorrAct}$ suggests potential corrective actions. This structured format provides actionable feedback that LLMs can leverage to improve fault localization and enable automated code repair.
4.2.2 Explicit Type Annotations
Type ambiguity commonly leads to erroneous code generation, particularly when dealing with closely related geometric entities such as Wire and Edge. ECIP mitigates this issue by enforcing explicit type annotations through a disciplined naming convention for functions, following the pattern:
$\mathtt{ActionName\_rReturnType},$ (9)
@@ 4.2.2 Explicit Type Annotations
where $\mathtt{ActionName}$ denotes the CAD operation, $\mathtt{r}$ is a fixed prefix indicating "returns", and $\mathtt{ReturnType}$ explicitly declares the type of the returned object. This ensures that even without inspecting documentation, the function signature itself conveys type information, reducing ambiguity during LLM inference. Each function is also accompanied by comprehensive, type-annotated documentation. This explicit type information reduces ambiguity and improves code generation accuracy by LLMs.
4.2.3 Self-Evolving Composite Operations
To facilitate knowledge accumulation and abstraction, ECIP supports the definition of composite operations—higher-level constructs formed by encapsulating sequences of atomic operations. These composites are saved as new operations within the language, enabling agents to invoke them directly in subsequent tasks.
This self-evolving mechanism effectively expands the language’s set of available operations, allowing reuse of complex modeling patterns such as screws, flanges, or other parametric features. Empirically, this contributes to improved one-shot generation success rates by allowing LLM agents to leverage accumulated domain knowledge efficiently.
4.3 Knowledge Base for Code Framework
LLMs possess strong generative and reasoning abilities but often struggle with the fine-grained semantic requirements of CAD APIs, such as precise parameter formats, geometric constraints, and operation-specific return types. These limitations frequently lead to misinterpreted command signatures or hallucinated inputs, resulting in unstable code generation.
To provide explicit, domain-grounded guidance during modeling, we introduce a structured knowledge base $\mathcal{K}$ coupled with retrieval-augmented generation (RAG). Retrieved function annotations and example cases supply accurate API semantics that complement the LLM’s implicit priors, substantially improving generation reliability. The knowledge base $\mathcal{K}$ consists of two key components:
@@ 4.3 Knowledge Base for Code Framework
  • Function Annotations ($\mathtt{Anno}$): semantic descriptions of API functions, including parameter formats, return types, and usage notes;
  • Case Examples ($\mathtt{Case}$): Initially composed of manually selected examples, and gradually expanded during use.
@@
To automate the construction of the knowledge base $\mathcal{K}$, we design a rule-based pipeline that extracts and organizes information directly from the source code. Specifically, we collect function signatures, parameter types, and return types from Python annotations and type hints. These are then aligned with the corresponding natural language descriptions provided in the function-level docstrings.
@@
5 Experiments and Comparison
This section details the implementation of CADDesigner and the comprehensive experimental evaluation conducted to assess its performance in CAD modeling code generation.
5.1 Implementation Details
We evaluate our CAD modeling code generation algorithm on a workstation with an 8-core Intel CPU under Python 3.12. The system is implemented on a customized agent framework that leverages retrieval-augmented generation (RAG) via RAGFlow. For embedding-based retrieval, we utilize the embedding model bge-large-zh-v1.5 to encode knowledge base documents. The retrieval employs a hybrid similarity scoring method, where vector similarity and keyword-based similarity are combined with weights of 0.6 and 0.4, respectively. The top 3 most relevant documents (top-k=3) are retrieved for downstream generation. For LLM services, the primary agent and the code generation tool ($T_{2}$) employ Claude-4-Sonnet, while requirement refinement ($T_{1}$) and both visual feedback sub-tools (visual question generation and visual feedback generation) are based on Gemini-2.5-Flash. Inference uses a temperature of 0.7 and nucleus sampling (top-p) of 0.9 to balance diversity and coherence. Detailed prompt designs, semantic chunking strategies, and other related details are described in the Appendix.
5.2 Dataset
Our experiments utilize the large-scale DeepCAD dataset [wu2021deepcad], which contains approximately 178K parametric CAD models based on the sketch-and-extrude representation. Following the approach in SkexGen [xu2022skexgen], we first remove duplicate models to reduce redundancy. Subsequently, we apply a filtering strategy inspired by HNC-CAD [xu2023hierarchical], retaining only those CAD models where the number of solids does not exceed 5, the total number of loops in all sketches is at most 20, the total number of curves in all loops is at most 60, and the total number of commands in the sketch-and-extrude sequence does not exceed 200. This two-stage filtering result in a refined dataset more suitable for downstream modeling tasks. From the filtered training set, we randomly select 1K models to initialize the RAG knowledge base. Deduplication ensures these models do not overlap with those in the test set. Additionally, a conversion script has been developed to transform the parametric command sequences of each model into ECIP code snippets, which are included in the RAG knowledge base as $\mathtt{Case}$. For evaluation, we randomly sample 200 and 1K models from the filtered test set for ablation studies and comparison experiments, respectively. For each sampled model, a single representative image is rendered from a fixed camera viewpoint to provide visual input in the ablation experiments.
5.3 Evaluation Metrics
To assess the geometric fidelity of the generated CAD models, we adopt several standard shape similarity metrics: Intersection over Union (IoU), Chamfer Distance (CD), and Hausdorff Distance (HD). These metrics evaluate how closely the generated geometry aligns with the ground-truth models from multiple perspectives. In addition, we introduce three process-oriented metrics to analyze modeling stability and reliability under ablation settings: Pass@1 denotes the percentage of cases where valid code is generated on the first attempt, reflecting generation accuracy. AVG Re (Average Retry) captures the mean number of retries required during the iterative correction loop, indicating convergence efficiency. SUC denotes the proportion of cases where a syntactically valid CAD model is successfully generated, i.e., models that can be executed and rendered without errors. Note that a syntactically valid model is not necessarily semantically aligned with the input description. Detailed mathematical definitions and metric computation protocols are provided in Appendix.
5.4 Comprehensive Ablation Study
To comprehensively verify the design choices and the effectiveness of different components in our CADDesigner agent, we conducted two groups of ablation studies on the DeepCAD dataset: one focusing on the ECIP paradigm, and another evaluating the impact of CAD tools as well as input modalities.
5.4.1 Effect of ECIP Paradigm
@@
To illustrate the benefits of certain design choices in ECIP, such as the error handling mechanism that is friendly to LLMs and the overall architecture of the ECIP system itself, we conducted an ablation study on a subset of 200 models randomly sampled from the filtered test set described in Section 5.2. Specifically, we evaluated four variants on the same dataset: CadQuery, ECIP, ECIP without the detailed error system, and ECIP without the return type annotations. The results are presented in Table 1.
@@
Table 1: Ablation study of ECIP components and comparative evaluation with CadQuery on 200 test models.
Pass@1↑ AVG Re↓ SUC↑
ECIP 0.46 1.85 100%
ECIP w/o Err 0.42 2.60 79.5%
ECIP w/o Type 0.31 2.30 90.5%
CadQuery 0.17 3.95 71.5%
@@
The results show that ECIP achieves the best overall performance, clearly outperforming CadQuery in all metrics. When the error handling mechanism is removed (ECIP w/o Err), the Pass@1 decreases slightly from 0.46 to 0.42, while the AVG Re increases from 1.85 to 2.60, and SUC drops to $79.5\%$. This suggests that although the agent can still generate correct code initially, the lack of detailed feedback slows down convergence during iterative correction and reduces the overall success rate. In contrast, removing type annotations (ECIP w/o Type) causes Pass@1 to drop significantly from 0.46 to 0.31, indicating that the initial code generation is more prone to semantic errors without explicit type guidance. The AVG Re (2.30) is higher than full ECIP, but still lower than ECIP without error handling, and SUC also drops to $90.5\%$. This shows that type annotations help encode critical modeling intent into the prompt, yet some type errors can still be corrected through retries. Overall, we observe a clear division of roles: error handling helps the agent quickly recover from mistakes during iterations, and type annotations improve first-attempt accuracy by reducing semantic ambiguity.
5.4.2 Effect of CAD Tools and Input Modalities
To further analyze the individual contributions of the key components in CADDesigner, we conducted an additional ablation study. We evaluated seven variants, labeled A to G, which respectively correspond to: A) removal of function annotations $\mathtt{Anno}$ from the knowledge base, B) removal of basic model cases $\mathtt{Case}$, C) disabling the requirement refinement tool $T_{1}$, D) disabling the visual feedback tool $T_{4}$, E) the full CADDesigner pipeline with text-only input, F) full pipeline with image-only input, and G) full pipeline with combined text and image input. Except for variants F and G, all others use text input exclusively.
Table 2: Ablation study of CADDesigner components and input modalities on 200 test models.
IoU↑ CD↓ HD↓
A (w/o $\mathtt{Anno}$)
B (w/o $\mathtt{Case}$) 0.2643 0.1320 0.4918
C (w/o $T_{1}$) 0.2713 0.1344 0.4764
D (w/o $T_{4}$) 0.2871 0.1160 0.4142
E (Text only) 0.3292 0.1148 0.3758
F (Image only) 0.3720 0.1028 0.3606
G (Text + Image) 0.3992 0.0826 0.2606
Table 2 summarizes the quantitative results across multiple evaluation metrics. In Model A, removing semantic API annotations ($\mathtt{Anno}$) prevents the agent from reliably identifying available operations and their usage, leading to complete failure in generating valid models. This indicates that structured function-level annotations play a critical role in grounding agent-based CAD code generation and guiding convergence toward correct solutions.
Models B, C, and D disable key system components—case examples ($\mathtt{Case}$), requirement refinement ($T_{1}$), and visual feedback ($T_{4}$), respectively. Compared to the text-only variant (Model E), all three models exhibit consistent performance degradation across metrics (IoU: 0.2643, 0.2713, 0.2871 vs. 0.3292), suggesting that each component contributes positively to robust and accurate model generation. Specifically, removing structured case examples and requirement refinement increases semantic ambiguity during early-stage generation, while disabling visual feedback primarily impairs iterative correction, as reflected by higher CD and HD values.
Switching the input modality from text-only (Model E) to image-only (Model F) yields a notable performance improvement, suggesting that abstract textual descriptions alone are often insufficient to fully specify geometric details, whereas images provide richer spatial cues for CAD modeling. Model G, which combines both text and image as input, achieves the best results across all metrics, illustrating that textual information can effectively complement visual context by resolving ambiguities difficult to infer from images alone.
Overall, these results indicate that structured API knowledge, coordinated tool usage, and multimodal input jointly contribute to high-fidelity and robust CAD model generation.
5.5 Comparison Methods